OpenStudy (anonymous):

What is the equation of a circle with center (2,7) and radius 4?

OpenStudy (anonymous):

\[ p : f(x,y)=(x-2)^2+(y-7)^2=16 \] General \[ f(x,y)=(x-u)^2+(y-v)^2=r^2\] where (u,v) is the center of the circle.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - h)^2 + (y - k)^2 = r^2}\) \((h,k)\) is the center, and \(r\) is the radius.

OpenStudy (anonymous):

a circle with centre at (a,b) and radius r has the equation: \[(x-a)^{2}+(y-b)^{2}=r ^{2}\]

OpenStudy (anonymous):

Thank you all, needed this extra help

Parth (parthkohli):

You're welcome!